Can you help me write these out in an algebraic expression?
1. Seven less than twice a number equals 19.

2.Eight decreased by 3 times a number equals 2.

3. The sum of two times a number and 5 is 11.


Is this right?
x/8 + -1/2 = 6

My answer: x=44
1
Expert's answer
2011-08-23T04:21:21-0400
1. 2x - 7 = 19 x = 13
2. 8 - 3x = 2 x = 2
3. 2x + 5 = 11 x = 3

x/8 + -1/2 = 6
x/8 - 0.5 = 6
x/8 = 6.5
x = 6.5* 8 = 52